# Optimizing against Infeasible Inclusions from Data for Semantic Segmentation through Morphology

### v4

@@ Abstract
State-of-the-art semantic segmentation models are typically optimized in a data-driven fashion, minimizing solely per-pixel or per-segment classification objectives on their training data. This purely data-driven paradigm often leads to absurd segmentations, especially when the domain of input images is shifted from the one encountered during training. For instance, state-of-the-art models may assign the label “road” to a segment that is included by another segment that is respectively labeled as “sky”. However, the ground truth of the existing dataset at hand dictates that such inclusion is not feasible. Our method, Infeasible Semantic Inclusions (InSeIn), first extracts explicit inclusion constraints that govern spatial class relations from the semantic segmentation training set at hand in an offline, data-driven fashion, and then enforces a morphological yet differentiable loss that penalizes violations of these constraints during training to promote prediction feasibility. InSeIn is a light-weight plug-and-play method, constitutes a novel step towards minimizing infeasible semantic inclusions in the predictions of learned segmentation models, and yields consistent and significant performance improvements over diverse state-of-the-art networks across the ADE20K, Cityscapes, and ACDC datasets. https://github.com/SHAMIK-97/PhyFea_on_Mask2Former/tree/main.
@@ 1 Introduction
Semantic segmentation is a fundamental task in computer vision and enables many downstream applications. The last decade has witnessed dramatic advances in this dense prediction task, driven by the introduction of end-to-end learned network architectures for solving it [30, 34, 5, 6, 47, 7, 20, 41, 50, 9, 48] and the development of ever stronger backbones coming from image classification [37, 21, 42, 14, 29, 49]. Indeed, designing a suitable backbone for dense semantic prediction that balances global context aggregation with preservation of fine local details has been a primary research direction in semantic segmentation. However, merely optimizing the backbone architecture neglects the structured prediction character of the problem. In particular, the simple optimization of cross-entropy loss terms even by state-of-the-art network architectures [41, 9] essentially poses semantic segmentation as a set of independent pixel-level or- at best—segment-level prediction problems and only loosely promotes regularity of the complete image-level outputs via the shared computation in the network across neighboring pixels. Very few deep-learning-based works [26] have focused on designing objectives that originate from the interaction between different pixels in the output predictions, and those that do focus only on local neighborhoods rather than large image regions. To the best of our knowledge, no previous method in semantic segmentation with deep learning has considered high-level pairwise objectives at the coarse level of semantic segments.
Aiming at promoting conformity of semantic segmentation models to high-level constraints, yet adhering to the taxonomy of the dataset at hand, we present a new method named Infeasible Semantic Inclusions (InSeIn). Our method has been empirically motivated by the observation that various state-of-the-art semantic segmentation networks [44, 9, 48, 1] make errors in their predictions, which are misclassifications or false positives, as illustrated in Fig. 1. In particular, based on the Cityscapes [11] taxonomy, a road segment cannot be entirely included in/surrounded by a sky segment, and the same holds for traffic sign and bus respectively. We term such segment inclusions infeasible inclusions and use them as an exemplary type of constraint that can provide a segmentation model with higher-level conformity based on the feasible spatial configurations of semantic segments in the ground truth.
@@
More specifically, the major contribution of InSeIn is a novel inclusion loss, which is used as an additional objective besides standard cross-entropy for optimizing a generic segmentation network. Our inclusion loss is computed based on the softmax outputs of the employed network, similarly to cross-entropy, which allows its general application to the vast majority of deep-learning-based semantic segmentation models. The key idea of computing this loss is to identify “soft” infeasible inclusions, which occur as regions in the softmax score maps of infeasible inclusion class pairs in which the ranking of the two softmax values is inverted compared to their surroundings. We recognize that this simply amounts to a grayscale morphological area opening [16] and derive a differentiable, parameter-free, and efficient implementation of it, which is normally included in the forward and backward pass of standard backpropagation.
@@ 2 Related Work
Figure 2: Overview of InSeIn. Top left: the complete network architecture, where the standard cross-entropy loss $l_{\text{ce}}$ from the baseline network $\phi$ is added to the inclusion loss $l_{\text{inclusion}}$ computed by InSeIn. Bottom: The pipeline of InSeIn. For the softmax outputs $P_{\phi}(I)$ of the network and for each pair of classes $(c_{i},c_{j})\in\mathcal{C^{\prime}}$ that signifies an infeasible inclusion where $c_{i}$ cannot include $c_{j}$, we take the difference of the respective softmax scores, $\text{P}_{\phi_{c_{j}}}(I)-\text{P}_{\phi_{c_{i}}}(I)$, and rectify it with a $\operatorname{ReLU}$. After concatenating all such $\mathcal{C^{\prime}}$ rectified difference maps channel-wise into $\hat{P}_{\phi}(I)$, we negate the latter and set its border pixels to 1. After that, an iterative operation for area opening is performed $T$ times on $\hat{P}_{\phi}(I)$. In each iteration, we perform max-pooling with a 3x3 kernel and a stride of 1, and the result is multiplied with $\hat{P}_{\phi}(I)$ element-wise. The final area-opened tensor channels differ from their counterparts in $\hat{P}_{\phi}(I)$ only across regions of incorrect inclusions. This element-wise difference is stored in $\mathbf{H}(I)$ and the $L_{1}$ norm of the latter, capturing both the spatial extent and the intensity of infeasible inclusions, constitutes the inclusion loss $l_{\text{inclusion}}$ employed in InSeIn.
@@ 3.1 Method Overview
$l_{\text{total}}=l_{\text{ce}}+\alpha{}l_{\text{inclusion}},0<\alpha<1.\vspace{-2mm}$ (1)
@@
1:$\mathcal{P}=\{(p,q):p\in\{1,\dots,C\},q\in\{1,\dots,C\}\setminus\{p\}\}$
2:$\mathcal{S}=\{I:\,I\in\mathbb{N}^{1,H,W}\}$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 3.3 Inclusion Loss
To compute our novel inclusion loss, we operate on the softmax maps $P_{\phi}(\mathbf{I})$ output by the segmentation network and index them channel-wise for each of the $|\mathcal{C}^{\prime}|$ infeasible inclusion pairs. Our core intuition is to identify in the pairs of softmax score maps that correspond to each infeasible inclusion case in $\mathcal{C}^{\prime}$ those regions across which class $c_{j}$—the includee—has a larger score than class $c_{i}$—the includer—and around which the opposite is true. These regions indicate “soft” infeasible inclusions in the softmax predictions, which should be penalized proportionally to the area they occupy and to the variable “intensity” of the inclusion, i.e. the magnitude of the difference between the softmax scores of classes $c_{i}$ and $c_{j}$ at each pixel of these regions. The computation of such a loss, which we term the inclusion loss, can be directly achieved in a differentiable formulation, amenable to standard backpropagation, via a grayscale morphological area opening of the aforementioned softmax difference maps. We implement this opening through an iterative algorithm, detailed in the following.
@@
where $N(p)$ denotes the structuring element of the dilation operator or, equivalently, the max-pooling kernel, which we set to an 8-connected neighborhood, i.e. a 3x3 kernel. We iterate (2) $T$ times, where $T=\max\{2,\min\{H,W\}/2\}$ and $H$ and $W$ are the height and width of the input image $\mathbf{I}$. Namely, we iterate for the minimal number of times that is required for initial positive values at image borders in $o^{(0)}_{(c_{i},c_{j})}$ to be able to reach any pixel in the interior. The final map $o^{(T)}_{(c_{i},c_{j})}$ is by construction 00 at all pixels that belong to infeasibly included regions. This property implies that
$\frac{\hat{P}_{\phi,(c_{i},c_{j})}(\mathbf{I})\odot o^{(T)}_{(c_{i},c_{j})}}{\max\left\{o^{(T)}_{(c_{i},c_{j})},\epsilon\right\}}\vspace{-2mm}$ (3)
@@ 4.2 Implementation Details
The softmax maps $P_{\phi}(\mathbf{I})$ input to InSeIn are first resized to $C\times 256\times 256$. The kernel size of the max-pooling used is $3\times 3$ with a stride of 1 and a padding of 1 to preserve spatial dimensions and to check for any connected components not sharing the border pixels. Random sampling on $|\mathcal{C^{\prime}}|$ is applied offline in each iteration to reduce the memory footprint. We report semantic segmentation performance using mean Intersection over Union (mIoU). The value of $\alpha$ is fixed to $10^{-9}$ for all examined datasets after a search in log-space in the interval $[10^{-10},10^{-5}]$ based on validation performance.
@@ 4.3 Comparison to The State of The Art
Tab. 1 compares InSeIn with a wide range of state-of-the-art methods for semantic segmentation on Cityscapes [11], ADE20K [52], and ACDC [36]. Overall, InSeIn achieves significant improvements in mIoU across all five network architectures on which we have implemented it and across all three datasets. On Cityscapes, we observe a consistent 0.50.50.50.50.80.80.80.8% improvement in mIoU with InSeIn on all architectures. On ADE20K, InSeIn achieves $1.2$% improvement over the baseline OCRNet [48] and $0.7$% over the baseline SegFormer-B4. On ACDC, InSeIn achieves $1.1$% improvement over the baseline OCRNet and $2.1$% over the baseline SegFormer-B4. Overall, improvements with InSeIn are consistent across all examined network architectures. This evidence shows that InSeIn can improve the segmentation performance of a general baseline network on which it is applied.
@@
Tab. 2 presents a comparison of InSeIn with a recent competing plug-and-play model, Beyond Pixels [22], on the Cityscapes dataset, implemented on two of our examined baseline networks, namely SegFormer and Mask2Former, for fully supervised segmentation. InSeIn consistently outperforms Beyond Pixels. Tab. 3 shows the class-wise comparison of IoU on the Cityscapes validation set for InSeIn and the corresponding baselines. We observe that InSeIn achieves a consistent improvement in IoU scores across most of the 19 classes. InSeIn with OCRNet [48], SegMAN [15], and SegFormer [44] show improvement in small challenging classes of the Cityscapes dataset. While we can observe an all-around and consistent improvement with Mask2Former [9] and OneFormer [25]. Tab. 4 shows the class-wise comparison of IoU on the ACDC test set for InSeIn and the corresponding baselines. We also observe that InSeIn achieves an overall improvement in IoU scores across the classes, especially over the SegFormer baseline. InSeIn achieves significant gains across all baselines for challenging small-instance classes such as motorcycle, bicycle, and rider.
@@
Qualitative results on ACDC.  In Fig. 6, in the first row, baseline prediction, wall segment infeasibly includes a sidewalk segment marked with green boxes. InSeIn successfully rectifies the infeasibility. The second row, baseline prediction, building segment infeasibly includes a wall segment marked with green boxes. InSeIn successfully predicts a wall label on that region. In both examples, red boxes highlight miscellaneous parts of the predictions where InSeIn improves upon baseline.
@@ 4.4 Evaluation of Prediction Feasibility
$\centering\text{mINF}=\frac{1}{|\mathcal{C}^{\prime}|}\sum_{(c_{i},c_{j})\in\mathcal{C}^{\prime}}\frac{f_{\text{inclusion}}(c_{i},c_{j})}{f_{\text{co-occur}}(c_{i},c_{j})}.\vspace{-2mm}\@add@centering$ (6)
@@
We additionally conduct an “oracle” experiment, by manually filtering with a conceptual process the original set of infeasible class pairs extracted from Cityscapes, which contains 180 pairs. After this filtering, the number of infeasible pairs becomes 120. The mIoU on the Cityscapes test set of the InSeIn model trained with this updated set of infeasible pairs based on a Mask2Former architecture is 83.9%, which is on a par with the original score presented in Tab. 1. This confirms that our data-driven constraint extraction already provides a meaningful set of constraints for moderate-size datasets such as Cityscapes, with minimal to no impact from potential spurious infeasibility constraints that are included to our loss due to the rare occurrence of the respective configurations in practice.

### v7

@@ Abstract
State-of-the-art semantic segmentation models are typically optimized in a data-driven fashion, minimizing solely per-pixel or per-segment classification objectives on their training data. This purely data-driven paradigm often leads to absurd segmentations, especially when the domain of input images is shifted from the one encountered during training. For instance, state-of-the-art models may assign the label “road” to a segment that is included by another segment that is respectively labeled as “sky”. However, the ground truth of the existing dataset at hand dictates that such inclusion is not feasible. Our method, Infeasible Semantic Inclusions (InSeIn), first extracts explicit inclusion constraints that govern spatial class relations from the semantic segmentation training set at hand in an offline, data-driven fashion, and then enforces a morphological yet differentiable loss that penalizes violations of these constraints during training to promote prediction feasibility. InSeIn is a light-weight plug-and-play method, constitutes a novel step towards minimizing infeasible semantic inclusions in the predictions of learned segmentation models, and yields consistent and significant performance improvements over diverse state-of-the-art networks across the ADE20K, Cityscapes, and ACDC datasets. Codebase will be made available. Code is available at https://github.com/SHAMIK-97/InSeIn
@@ 1 Introduction
Semantic segmentation is a fundamental task in computer vision and enables many downstream applications. The last decade has witnessed dramatic advances in this dense prediction task, driven by the introduction of end-to-end learned network architectures for solving it [30, 34, 5, 6, 49, 7, 20, 43, 52, 9, 50] and the development of ever stronger backbones coming from image classification [38, 21, 44, 14, 29, 51]. Indeed, designing a suitable backbone for dense semantic prediction that balances global context aggregation with preservation of fine local details has been a primary research direction in semantic segmentation. However, merely optimizing the backbone architecture neglects the structured prediction character of the problem. Since structured prediction means that pixel-labels are not independent but are spatially and semantically correlated. In particular, the simple optimization of cross-entropy loss terms even by state-of-the-art network architectures [43, 9] essentially poses semantic segmentation as a set of independent pixel-level or- at best—segment-level prediction problems and only loosely promotes regularity of the complete image-level outputs via the shared computation in the network across neighboring pixels. Very few deep-learning-based works [26] have focused on designing objectives that originate from the interaction between different pixels in the output predictions, and those that do focus only on local neighborhoods rather than large image regions. To the best of our knowledge, no previous method in semantic segmentation with deep learning has considered high-level pairwise objectives at the coarse level of semantic segments.
Aiming at promoting conformity of semantic segmentation models to high-level constraints, yet adhering to the taxonomy of the dataset at hand, we present a new method named Infeasible Semantic Inclusions (InSeIn). Our method has been empirically motivated by the observation that various state-of-the-art semantic segmentation networks [46, 9, 50, 1] make errors in their predictions, which are misclassifications or false positives, as illustrated in Fig. 1. In particular, based on the Cityscapes [11] taxonomy, a road segment cannot be entirely included in/surrounded by a sky segment, i.e. it cannot happen that the entire boundary of the road segment coincides with a given sky segment, and the same holds for traffic sign and bus respectively. We term such segment inclusions infeasible inclusions and use them as an exemplary type of constraint that can provide a segmentation model with higher-level conformity based on the feasible spatial configurations of semantic segments in the ground truth.
@@
More specifically, the major contribution of InSeIn is a novel inclusion loss, which is used as an additional objective besides standard cross-entropy for optimizing a generic segmentation network. Our inclusion loss is computed based on the softmax outputs of the employed network, similarly to cross-entropy, which allows its general application to the vast majority of deep-learning-based semantic segmentation models. The key idea of computing this loss is to identify “soft” infeasible inclusions, which occur as regions in the softmax score maps of infeasible inclusion class pairs in which the ranking of the two softmax values is inverted compared to their surroundings. We recognize that this simply amounts to a grayscale morphological area opening [16, 41] and derive a differentiable, parameter-free, and efficient implementation of it, which is normally included in the forward and backward pass of standard backpropagation.
@@ 2 Related Work
Figure 2: Overview of InSeIn. Top left: the complete network architecture, where the standard cross-entropy loss $l_{\text{ce}}$ from the baseline network $\phi$ is added to the inclusion loss $l_{\text{inclusion}}$ computed by InSeIn. cf. Sec. 3.1. Bottom: The pipeline of InSeIn, cf. Sec. $3.$ For the softmax outputs $P_{\phi}(I)$ of the network and for each pair of classes $(c_{i},c_{j})\in\mathcal{C^{\prime}}$ that signifies an infeasible inclusion where $c_{i}$ cannot include $c_{j}$, we take the difference of the respective softmax scores, $\text{P}_{\phi_{c_{j}}}(I)-\text{P}_{\phi_{c_{i}}}(I)$, and rectify it with a $\operatorname{ReLU}$. After concatenating all such $\mathcal{C^{\prime}}$ rectified difference maps channel-wise into $\hat{P}_{\phi}(I)$, we negate the latter and set its border pixels to 1. After that, an iterative operation for area opening is performed $T$ times on $\hat{P}_{\phi}(I)$. In each iteration, we perform max-pooling with a 3x3 kernel and a stride of 1, and the result is multiplied with $\hat{P}_{\phi}(I)$ element-wise. The final area-opened tensor channels differ from their counterparts in $\hat{P}_{\phi}(I)$ only across regions of incorrect inclusions. This element-wise difference is stored in $\mathbf{H}(I)$ and the $L_{1}$ norm of the latter, capturing both the spatial extent and the intensity of infeasible inclusions, constitutes the inclusion loss $l_{\text{inclusion}}$ employed in InSeIn.
@@ 3.1 Method Overview
$l_{\text{total}}=l_{\text{ce}}+\alpha{}l_{\text{inclusion}},0<\alpha<1.\vskip-5.69054pt$ (1)
@@
[…]
@@
12:     $\texttt{count\_freq}\leftarrow\texttt{count\_freq}+1$
@@
14:   component , noOFLabels $\leftarrow\textbf{C}(\texttt{mask\_2})$
@@
17:     comp_filled $\leftarrow\textbf{F}(\texttt{comp\_mask})$
18:     contour $\leftarrow\textbf{Cont}(\texttt{comp\_filled})$
19:     result $\leftarrow\texttt{contour}\land\texttt{mask\_1}$
@@
26:  if $\texttt{count}/\texttt{count\_freq}*100>2$ then
@@ 3.3 Inclusion Loss
To compute our novel inclusion loss, we operate on the softmax maps $P_{\phi}(\mathbf{I})$ output by the segmentation network and index them channel-wise for each of the $|\mathcal{C}^{\prime}|$ infeasible inclusion pairs. Our core intuition is to identify in the pairs of softmax score maps that correspond to each infeasible inclusion case in $\mathcal{C}^{\prime}$ those regions across which class $c_{j}$—the includee—has a larger score than class $c_{i}$—the includer—and around which the opposite is true. These regions indicate “soft” infeasible inclusions in the softmax predictions, which should be penalized proportionally to the area they occupy and to the variable “intensity” of the inclusion, i.e. the magnitude of the difference between the softmax scores of classes $c_{i}$ and $c_{j}$ at each pixel of these regions. The computation of such a loss, which we term the inclusion loss, can be directly achieved in a differentiable formulation, amenable to standard backpropagation, via a grayscale morphological [16] area opening of the aforementioned softmax difference maps. We implement this opening through an iterative algorithm, detailed in the following.
@@
where $N(p)$ denotes the structuring element of the dilation operator or, equivalently, the max-pooling kernel, which we set to an 8-connected neighborhood, i.e. a 3x3 kernel. We iterate (2) $T$ times, where $T=\max\{2,\min\{H,W\}/2\}$ and $H$ and $W$ are the height and width of the input image $\mathbf{I}$. Namely, we iterate for the minimal number of times that is required for initial positive values at image borders in $o^{(0)}_{(c_{i},c_{j})}$ to be able to reach any pixel in the interior. The final map $o^{(T)}_{(c_{i},c_{j})}$ is by construction 0 at all pixels that belong to infeasibly included regions. This property implies that
$\frac{\hat{P}_{\phi,(c_{i},c_{j})}(\mathbf{I})\odot o^{(T)}_{(c_{i},c_{j})}}{\max\left\{o^{(T)}_{(c_{i},c_{j})},\epsilon\right\}}\vskip-5.69054pt$ (3)
@@ 4.2 Implementation Details
The softmax maps $P_{\phi}(\mathbf{I})$ input to InSeIn are first resized to $C\times 256\times 256$. The kernel size of the max-pooling used is $3\times 3$ with a stride of 1 and a padding of 1 to preserve spatial dimensions and to check for any connected components not sharing the border pixels. Random sampling on $|\mathcal{C^{\prime}}|$ is applied offline in each iteration to reduce the memory footprint. We report semantic segmentation performance using mean Intersection over Union (mIoU). The value of $\alpha$ is fixed to $10^{-9}$ for all examined datasets after a search in log-space in the interval $10^{-10},10^{-5}]$ based on validation performance.
@@ 4.3 Comparison to The State of The Art
Tab. 1 compares InSeIn with a wide range of state-of-the-art methods for semantic segmentation on Cityscapes [11], ADE20K [54], and ACDC [36]. Overall, InSeIn achieves significant improvements in mIoU across all five network architectures on which we have implemented it and across all three datasets. On Cityscapes, we observe a consistent 0.50.50.80.8% improvement in mIoU with InSeIn on all architectures. We have also included a comparison of our method with another light-weight morphological loss clDice [37], with Mask2Former on Cityscapes. Our method beats the loss. On ADE20K, InSeIn achieves 1.21.2% improvement over the baseline OCRNet [50] and 0.70.7% over the baseline SegFormer-B4. On ACDC, InSeIn achieves 1.11.1% improvement over the baseline OCRNet and 2.12.1% over the baseline SegFormer-B4. Overall, improvements with InSeIn are consistent across all examined network architectures. This evidence shows that InSeIn can improve the segmentation performance of a general baseline network on which it is applied.
@@
Tab. 2 presents a comparison of InSeIn with a recent competing plug-and-play model, Beyond Pixels [22], on the Cityscapes dataset, implemented on two of our examined baseline networks, namely SegFormer and Mask2Former, for fully supervised segmentation. InSeIn consistently outperforms Beyond Pixels. Tab. 3 shows the class-wise comparison of IoU on the Cityscapes validation set for InSeIn and the corresponding baselines. We observe that InSeIn achieves a consistent improvement in IoU scores across most of the 19 classes. InSeIn with OCRNet [50], SegMAN [15], and SegFormer [46] show improvement in small challenging classes of the Cityscapes dataset. While we can observe an all-around and consistent improvement with Mask2Former [9] and OneFormer [25]. InSeIn also shows improvement over the baseline Mask2Former with another morphological loss, clDice [37]. Tab. 4 shows the class-wise comparison of IoU on the ACDC test set for InSeIn and the corresponding baselines. We also observe that InSeIn achieves an overall improvement in IoU scores across the classes, especially over the SegFormer baseline. InSeIn achieves significant gains across all baselines for challenging small-instance classes such as motorcycle, bicycle, and rider.
@@
Table 5: Comparison for domain shift Mask2Former and Mask2Former+InSeIn trained on Cityscapes [11] and evaluated on ACDC [36].
Method mIoU (%)
Mask2Former [9] 62.2
InSeIn w/ Mask2Former (ours) 63.6
@@
Qualitative results on ACDC.  In Fig. 6, in the first row, baseline prediction, wall segment infeasibly includes a sidewalk segment marked with green boxes.InSeIn successfully rectifies the infeasibility. The second row, baseline prediction, building segment infeasibly includes a wall segment marked with green boxes.InSeIn successfully predicts a wall label in that region. In both examples, red boxes highlight miscellaneous parts of the predictions where InSeIn improves upon baseline.
@@ 4.4 Evaluation of Prediction Feasibility
$\centering\text{mINF}=\frac{1}{|\mathcal{C}^{\prime}|}\sum_{(c_{i},c_{j})\in\mathcal{C}^{\prime}}\frac{f_{\text{inclusion}}(c_{i},c_{j})}{f_{\text{co-occur}}(c_{i},c_{j})}.\vskip-5.69054pt\@add@centering$ (6)
@@
Limitation and discussion.  We conduct an “oracle” experiment, by manually filtering with a conceptual process the original set of infeasible pairs extracted from Cityscapes, which contains 180 pairs. After this filtering, the number of infeasible pairs becomes 120. The mIoU on the Cityscapes test set of the InSeIn model trained with this updated set of infeasible pairs based on Mask2Former architecture is 83.9%, which is on a par with the original score presented in Tab. 1. This confirms that our data-driven constraint extraction already provides a meaningful set of constraints for moderate-size datasets like Cityscapes, with minimal to no impact from potential spurious infeasibility constraints that are included at our loss due to the rare occurrence of the respective configurations in practice. But like all data-driven architectures, our model is also dependent on the taxonomy of the dataset. So, our model is also subjected to the noise of the dataset.
Evaluation on Domain Shift  We evaluated our method with respect to domain shift in adverse weather conditions. We took the baseline Mask2Former [9] trained on Cityscapes [11] and evaluated on our ACDC [36] dataset and followed the same procedure with Mask2Former w/ InSeIn, whose results are shown in Tab 5. It shows our method has achieved significant improvement, thereby showing robustness to domain shift.